OpenStudy (dsappleby):

Write (9+6i)-(1+3i) as a complex number in standard form

OpenStudy (zacksibayan):

9+6i-1+3i =9-1+6i+3i =8+9i

OpenStudy (4everaddicted2anime):

First you need to distribute the negative sign into (1+3i) After doing this you should get (-1-3i) Now all that is left to do is (9+6i)+(-1-3i) Drop the parenthesis to get 9+6i-1-3i Combine like terms( the first pair of like terms are 9 and -1) ( the second pair of like terms are 6i and -3i) to get 8+3i